### v1

@@ Abstract
We develop an overlapping generations model where each agent observes a verifiable private signal about the state and, with positive probability, also receives signals disclosed by his predecessor. The agent then takes an action and decides which signals to pass on. Each agent’s action has a positive externality on his predecessor and his optimal action increases in his belief about the state. We show that as the communication friction vanishes, agents become increasingly selective in disclosing information. As the probability that messages reach the next generation approaches one, all signals except those with the highest likelihood ratio will be concealed in equilibrium.
@@ 1 Introduction
We study steady state equilibria of a doubly infinite time horizon game. The state of the world is constant over time and is either high or low. In each period, only one agent is active. He observes a private signal about the state and, with some probability, also receives the signals his immediate predecessor passed on to him. The agent then chooses both an effort level on the real line and a subset of the signals he possesses—including his private signal and any inherited signals—to pass on to his immediate successor. With some probability, communication succeeds and his successor observes this subset of signals. However, agents cannot directly observe others’ actions, whether and when communication failed, and the exact sequence of the disclosed signals.
@@
Each agent’s optimal effort strictly increases in his belief about the state and his action has a positive externality on his predecessor. Therefore, when deciding which subset of signals to disclose, each agent’s objective is to maximize his successor’s belief about the state.
@@
Due to the plethora of equilibria, some of which are driven by unreasonable off-path beliefs, [^1] For example, when there are three or more signal realizations, there may exist equilibria where signal realizations with the highest likelihood ratio are concealed while those with the second-highest likelihood ratio are revealed. Such equilibria can be sustained by off-path beliefs in which an agent who observes the disclosure of any signal with the highest likelihood ratio infers that many signals with low likelihood ratios are concealed. we restrict attention to equilibria that satisfy at least one of the following two refinements, both are satisfied in all equilibria of Dye (1985)’s model. The first refinement requires that under agents’ equilibrium strategy, they will disclose all available signals with likelihood ratios above a threshold and will conceal all other signals, where the threshold can depend on agents’ private signals and signals they inherited. The second refinement requires agents’ beliefs to be monotone, in the sense that their belief about the state increases when their predecessor reveals one fewer lower‑likelihood‑ratio signal and reveals one more higher‑likelihood‑ratio signal.
@@
Theorem 1 shows that when communication fails with low probability, disclosure becomes extremely selective in the sense that in all strict equilibria that satisfy at least one of our two refinements, all signals other than the ones with the highest likelihood ratio are concealed.
@@
To better understand how disclosure selectiveness depends on communication frictions, Theorem 2 allows for arbitrary communication failure rates and focuses on equilibria in which agents disclose a signal if and only if its likelihood exceeds a cutoff that is common across all information sets (i.e., constant-threshold equilibrium). The theorem shows that, for every signal with a likelihood ratio greater than one, there exists a constant-threshold equilibrium in which the signal is revealed if and only if the communication failure rate exceeds a cutoff. This cutoff is strictly decreasing in the signal’s likelihood ratio. Consequently, as communication frictions vanish, disclosure becomes more selective, in the sense that a smaller set of signals is revealed.
@@
Our findings contrast with a well-known result in Dye (1985), which shows that a higher probability that the sender possesses evidence leads to less selective disclosure. This difference arises from the different forms of uncertainty agents face about others’ information structures. In our model, consider the sequence of realized signals following the last communication failure. This sequence consists of multiple good signals (those that are disclosed), interspersed with subsequences of bad signals (which may be of zero length). That is, there is one bad sequence preceding the first good signal, one between each pair of adjacent good signals, and one following the last good signal. Moreover, conditional on the state, the expected length of each bad subsequence is independent of the disclosed good signals. Hence, by disclosing another good signal, the agent also reveals the existence of another bad subsequence. This negative effect is absent in Dye (1985) as any disclosure proves that no information is hidden. As communication frictions vanish, the expected length of each bad subsequence increases while the likelihood ratio of each good signal remains unchanged. Disclosure becomes more selective since each disclosed good signal needs to have a higher likelihood ratio in order to overcome the stronger negative effect of another bad subsequence.
We illustrate our model with an application to the “belief in a just world” framework of Benabou and Tirole (2006). We add intergenerational selective disclosure to their environment in which young agents choose effort based on beliefs about the returns to effort, while the old agents benefit from young agents’ effort and thus have incentives to shape their beliefs by selectively transmitting verifiable anecdotes. Different steady-state disclosure conventions (e.g., transmitting no signals versus transmitting only the strongest evidence of high returns to effort) generate distinct long-run belief distributions and effort outcomes, naturally echoing the “American Dream” versus “Europe” comparison.
@@
Our work contributes to the literature on disclosure games where the sender has unknown amount of signals. [^2] Di Tillio, Ottaviani, and Sørensen (2021) study an agent who observes the realizations of the $k$ $(\leq n)$ highest draws among $n$ conditionally i.i.d. signals, which contrasts to our model where the number of signals disclosed is endogenous. In Bardhi and Bobkova (2023), each sender decides whether to disclose his signal without observing its realization. In Antić and Chakraborty (2025), the state is $n$-dimensional and the sender selects which of $k$ $(\leq n)$ realized values of those dimensions to disclose (typically not the highest ones). Onuchic and Ramos (2025) consider multiple senders who decide whether to disclose a multi-dimensional state via a collective decision rule. The models of Shin (1994, 2003) and Dziuda (2011) focus on signals with only two realizations, which cannot answer questions related to the selectiveness of disclosure. Gao (2025) studies a model where the sender has a continuum of signals and focuses on equilibria satisfying the truth-leaning refinement of Hart, Kremer, and Perry (2017), which conflicts with our refinements, as we explain in Section 3. She shows that the sender replicates the signal distribution under a more favorable state by disclosing some unfavorable signals. In contrast, we find that agents disclose only the most favorable signals as communication frictions vanish.
Rappoport (2025) focuses on the receiver’s optimal equilibrium and shows that when the sender is expected to have more information, the receiver will take actions that are worse for the sender in every state. Our analysis instead focuses on the selectiveness of the sender’s disclosure behavior rather than the receiver’s action choices. Moreover, we study threshold equilibria and equilibria with monotone beliefs, which may not include the receiver’s optimal equilibrium.
@@
A growing literature studies selective disclosure in dynamic settings. [^3] In static disclosure games where the number of signals available to the sender is drawn from a geometric distribution, whether the conclusion of our Theorem 1 continues to hold remains an open question. We will explain the challenges in analyzing the static disclosure game in the conclusion section. In Guttman, Kremer, and Skrzypacz (2014), a sender may receive up to two signals over two periods and chooses both what to disclose and when. In Felgenhauer and Schulte (2014), Felgenhauer and Loerke (2017), Lou (2023), Arieli and Stewart (2025), and Dai, Fudenberg, and Pei (2026), a sender conducts a sequence of experiments before selecting a subset of outcomes to disclose to a receiver. In Gieczewski (2022) and Squintani (2025), a subset of agents observe the state and decide whether to disclose it to their neighbors. Pei (2025) studies repeated games with random matching where players can voluntarily disclose signals about their past actions. To the best of our knowledge, none of these existing works examine the selectiveness of players’ disclosure behavior.
@@
Our work also contributes to the study of overlapping generations models starting from Samuelson (1958). While many papers in this literature (Bhaskar, 1998; Acemoglu and Jackson, 2015) focus on the observability of agents’ actions, we study settings where actions are not observed, so that agents can influence their successors’ behaviors only by disclosing verifiable signals. Niehaus (2011) studies the sharing of skills among a sequence of agents, which contrasts to our model where agents share verifable signals about some payoff-relevant state.
More closely related is the work of Anderlini, Gerardi, and Lagunoff (2012) that also analyzes an overlapping generations model with strategic communication and payoff externalities, and examines whether agents learn the state in the long run. Different from our model, communication in their model takes the form of cheap talk rather than the disclosure of hard evidence.
2 Model
@@
Consider a game with a doubly infinite time horizon $t\in\mathbb{Z}\equiv\{...-1,0,1,$...\}. Agent $t$ is the only active player in period $t$. He observes a private signal $s^{t}\in S$ about the state $\theta\in\{\underline{\theta},\overline{\theta}\}$, and possibly, some signals disclosed to him by agent t1t-1. Then he chooses an action $a^{t}\in\mathbb{R}$ and a set of signals to reveal to agent $t+1$. We assume that $\theta$ is constant over time, the set of signal realizations $S\equiv\{s_{1},s_{2},...,s_{l}\}$ is finite with $l\geq 2$, and that no signal perfectly reveals any state. Let $\pi_{0}\in(0,1)$ denote the prior probability that $\theta=\overline{\theta}$. Let $f_{\theta,i}$ denote the probability of $s_{i}$ conditional on state $\theta$ and let $\mathbf{f}\equiv\{f_{\overline{\theta},i},f_{\underline{\theta},i}\}_{i=1}^{l}$. To simplify notation, we write $\overline{f}_{i}$ instead of $f_{\overline{\theta},i}$ and $\underline{f}_{i}$ instead of $f_{\underline{\theta},i}$. Let $L_{i}\equiv\overline{f_{i}}/\underline{f_{i}}$ denote the likelihood ratio of signal $s_{i}$. We assume that $L_{i}\neq L_{j}$ for every $i\neq j$ and we index signals so that $L_{i}$ is strictly decreasing in $i$:
@@
Agent $t$’s payoff is $u(\theta,a^{t})+v(\theta,a^{t+1})$, which depends on the state $\theta$, his action, and his immediate successor’s action. We assume that (i) $v$ is strictly increasing in $a^{t+1}$, [^5] Similar results can be derived in the case where $v$ is strictly decreasing in $a^{t+1}$. (ii) $u$ is strictly concave in $a^{t}$, and that (iii) $\arg\max_{a^{t}\in\mathbb{R}}\{\pi u(\overline{\theta},a^{t})+(1-\pi)u(\underline{\theta},a^{t})\}$ is strictly increasing in $\pi\in[0,1]$. A concrete example is a setting similar to the one in Benabou and Tirole (2006): Each agent $t$ chooses effort $a^{t}\in\mathbb{R}_{+}$ at a strictly increasing and convex cost $c(a^{t})$. He produces output $\theta a^{t}+(1-\theta)\varepsilon^{t}$, where $\varepsilon^{t}$ stands for a random shock that is independent of effort and $\theta$ stands for the marginal product of effort that is either high ($\theta=\overline{\theta}$) or low ($\theta=\underline{\theta}$) with $0<\underline{\theta}<\overline{\theta}<1$. As in Samuelson (1958), he consumes an exogenous fraction of his own output and transfers the rest to his immediate predecessor through some social security system.
@@
The information agent t1t-1 reveals to agent $t$ is characterized by $\mathbf{h}\equiv(h_{1},h_{2},...h_{l})\in H\equiv\mathbb{N}^{l}$ where $h_{i}$ stands for the number of signal $s_{i}$ revealed. With probability $\delta\in(0,1)$, communication succeeds, in which case agent $t$ observes the number of each signal revealed by agent t1t-1 but not the exact order of these signals, i.e., agent $t$’s history is $\mathbf{h}$. With probability $1-\delta$, communication fails in which case agent $t$’s history is $\mathbf{0}\equiv(0,...,0)$. Agents observe their histories and private signals but cannot directly observe the calendar times at which communication fails. We interpret $1-\delta$ as a communication friction and refer to $h_{1}+...+h_{n}$ as the length of history $(h_{1},...,h_{n})$.
@@
We assume that signals are verifiable so that they can be concealed but not falsified (Grossman, 1981; Milgrom, 1981; Dye, 1985). Formally, let $\mathbf{1}_{i}\in\mathbb{N}^{l}$ denote a vector where the $i$th entry is $1$ and all other entries are 0. After observing history $\mathbf{h}$ and receiving private signal $s_{i}$, agent $t$ can only disclose vectors that are no more than $\mathbf{h}+\mathbf{1}_{i}$. Hence, an agent can only reveal a subset of the signals he possesses, which include his private signal and the signals he inherited (if any).
@@
A pure strategy $\sigma:H\times S\rightarrow\mathbb{R}\times H$ maps an agent’s history and private signal to his action and the information he discloses to his immediate successor, subject to the feasibility constraint. Our solution concept is steady state Bayesian equilibrium (or equilibrium for short), which extends the notion of steady state equilibrium in Clark, Fudenberg, and Wolitzky (2021) and Pei (2025) to incomplete information games. An equilibrium of our game consists of (i) a pure strategy $\sigma$, (ii) two distributions $\underline{\mu},\overline{\mu}\in\Delta(H)$ over histories, and (iii) a belief map $\pi:H\rightarrow[0,1]$ which represents the probability that an agent assigns to state $\overline{\theta}$ after observing his history but before observing his own private signal. These three components satisfy (i) when all agents use strategy $\sigma$, there exists a steady state in which agents’ histories are distributed according to $\underline{\mu}$ conditional on $\theta=\underline{\theta}$ and are distributed according to $\overline{\mu}$ conditional on $\theta=\overline{\theta}$, (ii) $\pi(\mathbf{h})$ is derived via Bayes rule from $(\underline{\mu},\overline{\mu})$ at every $\mathbf{h}$ that occurs with positive probability, and that (iii) $\sigma$ maximizes an agent’s expected payoff when all agents’ beliefs after observing their histories are given by the mapping $\pi$. A strict equilibrium is one in which the strategy $\sigma$ is strictly optimal for the agents at each history-signal pair $(\mathbf{h},s)\in H\times S$.

### v2

@@ Abstract
We develop an overlapping generations model where each agent observes a verifiable private signal about the state and, with positive probability, also receives signals disclosed by his predecessor. The agent then takes an action and decides which signals to pass on. Each agent’s action has a positive externality on his predecessor and his optimal action increases in his belief about the state. We show that as the probability that messages reach the next generation approaches one, agents become increasingly selective in disclosing information. In the limit, all signals except for the most favorable ones will be concealed.
@@ 1 Introduction
We study steady state equilibria of a doubly infinite time horizon game. The state of the world is constant over time and is either high or low. In each period, only one agent is active. He observes a private signal about the state and, with some probability, receives the signals his immediate predecessor passed on to him. The agent then chooses both an effort level and a subset of the signals he possesses (including his private signal and any inherited signals) to pass on to his immediate successor. With some probability, communication succeeds and the successor observes this subset. However, agents cannot directly observe others’ actions, whether and when communication failed, or the exact sequence of the disclosed signals.
@@
Each agent’s optimal effort strictly increases in his belief about the state and this effort generates a positive externality for his predecessor. Therefore, when choosing the subset of signals to disclose, each agent’s objective is to maximize his successor’s belief about the state.
@@
Due to the plethora of equilibria, some of which are driven by unreasonable off-path beliefs, [^1] For example, when there are three or more signal realizations, there may exist equilibria where signal realizations with the highest likelihood ratio are concealed while those with the second-highest likelihood ratio are revealed. Such equilibria can be sustained by off-path beliefs in which an agent who observes the disclosure of any signal with the highest likelihood ratio infers that many signals with low likelihood ratios are concealed. we restrict attention to equilibria that satisfy at least one of the following refinements. The first refinement requires that agents’ beliefs be monotone, in the sense that their belief about the state increases when their predecessor substitutes a lower‑likelihood‑ratio signal for a higher‑likelihood‑ratio one. The second refinement requires a threshold equilibrium, whereby agents disclose a signal if and only if its likelihood ratio exceeds a threshold that may depend on their private and inherited signals. Both refinements are satisfied by all equilibria of Dye (1985)’s model.
@@
Theorem 1 shows that when communication fails with low probability, disclosure becomes extremely selective. In particular, in all strict equilibria that satisfy at least one of our two refinements, agents conceal all signals except the ones with the highest likelihood ratio.
@@
To better understand how selectively agents disclose as communication frictions vary, Theorem 2 allows for arbitrary communication success rates. It focuses on constant-threshold equilibria, that is, threshold equilibria with a disclosure threshold that is constant across all information sets. The theorem shows that, for every signal with a likelihood ratio greater than one, there exists a constant-threshold equilibrium where the signal is disclosed if and only if the communication success rate is below a cutoff. This cutoff is strictly increasing in the signal’s likelihood ratio. Consequently, as communication frictions vanish, disclosure becomes more selective, in the sense that a smaller set of signals can be revealed in constant-threshold equilibria.
@@
Our findings contrast with Dye (1985), who shows that a higher probability of the sender possessing evidence leads to less selective disclosure. This difference arises from the different forms of uncertainty agents face about others’ information structures. To build intuition, consider the sequence of realized signals following the last communication failure in our model. This sequence consists of multiple good signals (those that are disclosed), interspersed with (sub)sequences of bad signals (which may be of zero length). That is, there is one bad sequence preceding the first good signal, one between each pair of adjacent good signals, and one following the last good signal. Conditional on the state, the expected length of each bad sequence is independent of the disclosed good signals. Hence, by disclosing another good signal, the agent also reveals the existence of another bad sequence. This negative inference is absent in Dye (1985), where any disclosure implies that no information is hidden. As communication frictions vanish, the expected length of each bad sequence increases while the likelihood ratio of each good signal remains unchanged. Disclosure therefore becomes more selective, since each disclosed signal must have a higher likelihood ratio in order to overcome the stronger negative effect of one more bad sequence.
@@
Our work contributes to the literature on disclosure games where the sender has an unknown amount of signals. [^2] Di Tillio, Ottaviani, and Sørensen (2021) study an agent who observes the realizations of the $k$ $(\leq n)$ highest draws among $n$ conditionally i.i.d. signals, which differs from our model where the number of signals disclosed is endogenous. In Bardhi and Bobkova (2023), each sender decides whether to disclose his signal without observing its realization. In Antić and Chakraborty (2025), the state is $n$-dimensional and the sender selects which of $k$ $(\leq n)$ realized values of those dimensions to disclose (typically not the highest ones). Onuchic and Ramos (2025) consider multiple senders who decide whether to disclose a multi-dimensional state via a collective decision rule. Most of the existing works on this topic study one-shot games. The models of Shin (1994, 2003) and Dziuda (2011) focus on binary signals, which are not well-suited to study the selectiveness of disclosure. Gao (2025) assumes that the sender has a continuum of signals and analyzes equilibria satisfying the truth-leaning refinement of Hart, Kremer, and Perry (2017), which, as we discuss in Section 3, conflicts with our refinements. [^3] Gieczewski and Titova (2024) propose a coalitional proofness refinement and provide conditions under which their refinement coincides with the truth-leaning refinement and the receiver-optimality refinement. She shows that the sender sometimes discloses bad signals in order to replicate the signal distribution under another state, which contrasts with our findings. [^4] Gao (2025) also shows that there exists a receiver-optimal equilibrium where the sender discloses only the most favorable signal, although there can also be other receiver-optimal equilibria. Rappoport (2025) focuses on receiver-optimal equilibria and shows that the receiver will take worse actions for the sender when the sender is expected to have more information. We instead examine the selectiveness of the sender’s disclosure behavior.
Unlike in one-shot disclosure games, the signals available to the senders in our overlapping generations game are endogenously determined in equilibrium. This leads to qualitatively different findings compared to one-shot games, as we describe in Section 5.1.
@@
A growing literature studies disclosure in other dynamic settings. In Guttman, Kremer, and Skrzypacz (2014), a sender may receive up to two signals over two periods and chooses both what to disclose and when. In Felgenhauer and Schulte (2014), Felgenhauer and Loerke (2017), Lou (2023), Arieli and Stewart (2025), and Dai, Fudenberg, and Pei (2026), a sender conducts a sequence of experiments before selecting a subset of outcomes to disclose to a receiver. In Gieczewski (2022) and Squintani (2025), some agents know the state and can disclose it to their neighbors. In Kremer, Schreiber, and Skrzypacz (2024), a potentially informed sender decides whether to disclose the current value of a random walk. Pei (2025) studies repeated games where players can voluntarily disclose signals about their past actions. Our paper complements this literature by examining how the selectiveness of disclosure depends on communication frictions.
@@
Our work also contributes to the study of overlapping generations models starting from Samuelson (1958). While many papers in this literature (Bhaskar, 1998; Acemoglu and Jackson, 2015) focus on the observability of agents’ actions, we study settings where actions are not observed, so that agents can influence their successors’ behaviors only by disclosing information. Anderlini, Gerardi, and Lagunoff (2012) analyze an overlapping generations model with strategic communication and payoff externalities. Communication in their model is cheap talk rather than the disclosure of hard evidence. They also focus on a different question – whether agents learn the state in the long run – while we study the selectiveness of disclosure.
Broadly speaking, our model is also related to the literature on social learning such as Banerjee (1992), Bikhchandani, Hirshleifer, and Welch (1992), and Smith and Sørensen (2000). [^5] Niehaus (2011) studies the sharing of complementary or substitutable skills among a sequence of agents, which contrasts with our model where agents share noisy signals about a payoff-relevant state of the world. Unlike those models, where there are no payoff externalities and agents learn from others’ actions in addition to their private signals, our model incorporates payoff externalities and agents learn from signals explicitly disclosed by others.
2 Baseline Model
@@
Consider a game with a doubly infinite time horizon $t\in\mathbb{Z}\equiv\{...-1,0,1,$...\}. Agent $t$ is the only active player in period $t$. He observes a private signal $s^{t}\in S$ about the state $\theta\in\{\underline{\theta},\overline{\theta}\}$, and with probability $\delta\in(0,1)$, some signals disclosed to him by agent t1t-1. Then he chooses an action $a^{t}\in\mathbb{R}$ and a set of signals to reveal to agent $t+1$. We assume that $\theta$ is constant over time, the set of signal realizations $S\equiv\{s_{1},s_{2},...,s_{l}\}$ is finite with $l\geq 2$, and that no signal reveals any state. Let $\pi_{0}\in(0,1)$ denote the prior probability that $\theta=\overline{\theta}$. Let $f_{\theta,i}$ denote the probability of $s_{i}$ conditional on state $\theta$. To simplify notation, we write $\overline{f}_{i}$ instead of $f_{\overline{\theta},i}$ and $\underline{f}_{i}$ instead of $f_{\underline{\theta},i}$. Let $\mathbf{f}\equiv\{\overline{f_{i}},\underline{f_{i}}\}_{i=1}^{l}$. Let $L_{i}\equiv\overline{f_{i}}/\underline{f_{i}}$ denote the likelihood ratio of signal $s_{i}$. We assume that $L_{i}\neq L_{j}$ for every $i\neq j$ and we index signals so that $L_{i}$ is strictly decreasing in $i$
@@
Agent $t$’s payoff is $u(\theta,a^{t})+v(\theta,a^{t+1})$, which depends on the state $\theta$, his action $a^{t}$, and his immediate successor’s action $a^{t+1}$. We assume that (i) $v$ is strictly increasing in $a^{t+1}$, (ii) $u$ is strictly concave in $a^{t}$, which implies that there is a unique $a^{*}(\pi)\in\mathbb{R}$ for every $\pi\in[0,1]$ that solves $\max_{a\in\mathbb{R}}\big\{\pi u(\overline{\theta},a)+(1-\pi)u(\underline{\theta},a)\big\}$, and that (iii) $a^{*}(\pi)$ is strictly increasing in $\pi$.
@@
The information agent t1t-1 reveals to agent $t$ is characterized by $\mathbf{h}\equiv(h_{1},h_{2},...h_{l})\in H\equiv\mathbb{N}^{l}$ where $h_{i}$ stands for the number of signal $s_{i}$ revealed. With probability $\delta\in(0,1)$, communication succeeds, in which case agent $t$ observes the number of each signal revealed by agent t1t-1 but not the exact order of these signals, i.e., agent $t$’s history is $\mathbf{h}$. With probability $1-\delta$, communication fails in which case agent $t$’s history is $\mathbf{0}\equiv(0,...,0)$. We interpret $1-\delta$ as a communication friction. We refer to $h_{1}+...+h_{l}$ as the length of history $(h_{1},...,h_{l})$. Each agent only observes his history and private signal, but cannot directly observe his predecessors’ actions, the calendar times at which communication fails, the exact sequence of disclosed signals, and so on.
@@
We assume that signals are verifiable so that they can be concealed but not falsified (Grossman, 1981; Milgrom, 1981; Dye, 1985). Formally, let $\mathbf{1}_{i}\in\mathbb{N}^{l}$ denote a vector where the $i$th entry is $1$ and all other entries are 0. After observing history $\mathbf{h}$ and receiving private signal $s_{i}$, agent $t$ can only disclose vectors that are no more than $\mathbf{h}+\mathbf{1}_{i}$. That is, he can only reveal a subset of the signals he possesses, which include his private signal and the signals he inherited (if any).
@@
A pure strategy $\sigma:H\times S\rightarrow\mathbb{R}\times H$ maps an agent’s history and private signal to his action and the information he discloses to his immediate successor, subject to the feasibility constraint. Our solution concept is steady state Bayesian equilibrium (or equilibrium for short), which extends the notion of steady state equilibrium in Clark, Fudenberg, and Wolitzky (2021) and Pei (2025) to incomplete information games. A (pure-strategy) equilibrium of our game consists of (i) a pure strategy $\sigma$, (ii) two distributions $\underline{\mu},\overline{\mu}\in\Delta(H)$ over histories, and (iii) a belief map $\pi:H\rightarrow[0,1]$ which represents the probability that an agent assigns to state $\overline{\theta}$ after observing his history but before observing his own private signal. These three components satisfy (i) when all agents use strategy $\sigma$, there exists a steady state in which agents’ histories are distributed according to $\underline{\mu}$ conditional on $\theta=\underline{\theta}$ and are distributed according to $\overline{\mu}$ conditional on $\theta=\overline{\theta}$, (ii) $\pi(\mathbf{h})$ is derived via Bayes rule from $(\underline{\mu},\overline{\mu})$ at every $\mathbf{h}$ that occurs with positive probability, and that (iii) $\sigma$ maximizes an agent’s expected payoff when all agents’ beliefs after observing their histories are given by the mapping $\pi$. [^7] As in Clark, Fudenberg, and Wolitzky (2021), our solution concept requires all agents using the same pure strategy, i.e., the same mapping $\sigma:H\times S\rightarrow\mathbb{R}\times H$. Because the distribution over histories is at a steady state in our equilibrium, whether an agent observes calendar time does not affect his belief about the state. A strict equilibrium is one in which the equilibrium strategy $\sigma$ is strictly optimal for the agents at each history-signal pair $(\mathbf{h},s)\in H\times S$.
The existence of strict equilibrium will be implied by Proposition 1. Lemma 1 shows that each pure strategy induces a unique steady state history distribution conditional on each state.
  Lemma 1.
For every pure strategy $\sigma$ and conditional on every $\theta\in\{\overline{\theta},\underline{\theta}\}$, there exists a unique steady state distribution over histories $\mu_{\theta}\in\Delta(H)$.
The proof is in Appendix A.1.
Remarks:
One interpretation of our model is paternalistic parenting in the spirit of Doepke and Zilibotti (2017). Children decide how to allocate time between schoolwork and tempting but unproductive activities (e.g., playing video games). The state $\theta$ captures the return to studying in terms of future well-being and is either high ($\theta=\overline{\theta}$) or low ($\theta=\underline{\theta}$). Children devote more time to studying when they believe the return is higher. Parents, motivated by paternalistic concern, prefer their children to allocate more time to studying than to playing video games. Children cannot observe whether their parents themselves studied hard when they were young. However, parents can shape their children’s beliefs by sharing anecdotes—some based on their own observations (i.e., their private signals) and others passed down from their own parents. [^8] This is referred to as authoritative parenting in Doepke and Zilibotti (2017) that “parents can mold their children’s preferences so as to align them with their own, for instance, by emphasizing the virtue of hard work…
Our model can also be applied to an organizational setting. New members (e.g., workers) choose their effort levels subject to a convex cost, while senior members (e.g., managers) tell stories regarding the productivity of effort. Because workers are more inclined to work hard when they perceive their effort as more productive, managers—who do not internalize the workers’ effort costs—strategically share stories to induce higher effort from the workers. Some of these stories are personal observations while others are inherited from previous generations. Workers cannot observe their managers’ effort in previous periods. Workers age into managerial roles after one period and choose a set of stories to pass on to the next generation of workers.
Our overlapping generations game is related to, albeit qualitatively different from, a one-shot disclosure game where the number of verifiable signals the sender observes is drawn from a geometric distribution with parameter $\delta$. We comment on the qualitative differences between the equilibria of our game and those in the one-shot disclosure game in Section 5.1.
In Sections 5.2 and 5.3, we discuss two extensions of our baseline model. This includes when agents face an upper bound on the number of signals disclosed, or when the communication success rate depends on the set of signals disclosed.